# Decentralized real-time iterations for distributed NMPC

### v2

@@ Abstract
I Introduction
@@
Distributed control concerns the operation and control of cyber-physical systems, e.g., energy systems [Venkat2008] or robot formations [vanParys2017]. Coupling in systems of systems can occur in the dynamics, via constraints, or through a common objective. A key challenge for the design of optimization-based schemes for interconnected and cyber-physical systems is to reconcile cooperation with the computational burden, i.e., real-time feasibility is a must in applications. On the far end of the spectrum, decentralized control schemes do not allow for cooperation but also do not require communication between subsystems [Siljak1991]. The opposite is centralized MPC, where the network of subsystems is considered as one, large-scale, dynamical system which is controlled by a single controller.
@@
Model Predictive Control (MPC), also known as receding-horizon optimal control, has seen tremendous industrial success catalyzed by the development of numerical schemes tailored to the dynamics and to the problem formulation. Of particular importance for the implementation of centralized Nonlinear MPC (NMPC) are Real-Time Iteration (RTI) schemes [Diehl2002, Zavala2009, Wolf2016, Zanelli2021, Kapernick2014, Darup2019]. Instead of solving the Optimal Control Problem (OCP) to full accuracy in each control step, an RTI scheme applies only one or a few iterations of an optimization method.
@@
On the other hand, Distributed MPC (DMPC) decomposes the numerical optimization among the subsystems [Scattolini2009, Muller2017]. DMPC design proceeds mainly along two dimensions: (i) the OCP formulation and (ii) the implementation of an optimization algorithm with the desired degree of decomposition or decentralization.
@@
With respect to (i), one may design either an individual OCP for each subsystem [Dunbar2007, Muller2012, Varutti2012], or a centralized OCP for the whole cyber-physical system [Conte2016, Bestler2019]. Schemes with individual OCPs for the subsystems require few communication rounds per control step, enjoy small communication footprints, and allow for fast sampling rates [Dunbar2007, Muller2012, Varutti2012]. However, they only allow for limited cooperation. Distributed approaches built upon centralized OCPs generally require multiple communication rounds per control step, but also allow for cooperation as the control tasks of all subsystems are encoded in the centralized OCP. We therefore refer to the latter approach as cooperative DMPC in this article.
@@
With respect to (ii), numerous optimization algorithms have been proposed to solve centralized OCPs online [Stewart2011, Summers2012, Hours2016, Stomberg2022]. A distinction can be made between distributed and decentralized optimization methods. Distributed optimization splits most computations between the subsystems and there exists a central entity which coordinates the subsystems [Bertsekas1989]. Decentralized optimization only requires neighbor-to-neighbor communication without a coordinator [Nedic2018].
@@
Ultimately, cooperative DMPC aims to combine the high performance of centralized MPC with the favorable communication structure of decentralized optimization. Real-time requirements dictate that only a finite number of optimizer iterations can be executed in each control step, which limits performance and must be addressed in the stability analysis. While RTI schemes have been pivotal in bringing NMPC to industrial applications, distributed counterparts for similar results are not available. For linear systems, DMPC-specific OCP designs are presented in [Conte2016, Darivianakis2019] and stability under inexact optimization is analyzed in [Kohler2019, Giselsson2014]. For nonlinear systems, cooperative DMPC schemes with stability guarantees are presented in [Stewart2011, Bestler2019]. In both articles, a centralized OCP with a terminal penalty is formulated such that the OCP value function serves as a candidate Lyapunov function for the closed-loop system. Stability is ensured in two different ways: Either a feasible-side convergent optimization method is employed and stability follows from standard arguments [Stewart2011]. The drawback of this approach is that the optimization method requires a feasible initialization, which is difficult to implement in practice, and that each subsystem needs access to the dynamics of all subsystems in the network. The second option for guaranteeing stability is to solve the OCP with the Alternating Direction Method of Multipliers (ADMM) until a tailored ADMM stopping criterion is met [Bestler2019]. This approach presumes that ADMM converges linearly, i.e., sufficiently fast, to the OCP minimizer. However, such ADMM convergence guarantees for problems with non-convex constraint sets are, to the best of the authors’ knowledge, yet unavailable. Moreover, we note that the existing ADMM convergence guarantees for non-convex constraint sets [Themelis2020, Wang2019] so far do not allow for decentralized implementations.
@@
Consequently, the existing stability guarantees of cooperative DMPC for coupled nonlinear systems either require feasible initialization and global model knowledge in all subsystems or they rely on rather strong assumptions on the achieved optimizer convergence. To overcome these limitations, we propose a novel decentralized RTI scheme for distributed NMPC. Our scheme builds on a bi-level decentralized Sequential Quadratic Programming (dSQP) scheme [Stomberg2022a]. On the outer level, the method uses an inequality-constrained SQP scheme, which leads to partially separable convex Quadratic Programs (QPs) to be solved in each SQP step. On the inner level, these QPs are solved with ADMM which is guaranteed to converge and can be implemented in decentralized fashion, i.e., it does not require a central coordinator. Our previous conference paper [Stomberg2022a] presents an earlier dSQP version including a stopping criterion for ADMM, but it does not consider real-time control applications and stability analysis. The idea of executing only few iterations of a tailored optimization method to enable distributed NMPC in real-time is also used in [Hours2016]. Therein, an augmented Lagrangian-based decomposition scheme is proposed for non-convex OCPs and suboptimality bounds for the optimizer solutions are derived, if changes in the system state between subsequent NMPC steps are small. While stability of the dynamical system is not formally discussed in [Hours2016], the approach shares commonalities to the scheme developed in this article and we later give a more detailed comparison in Remark 8 in Section V.
@@
In the present paper, we explore the theoretical foundation of decentralized real-time iterations for DMPC via dSQP. The real-time feasibility of our approach has successfully been validated in experiments with mobile robots and on embedded hardware [Stomberg2023, Stomberg2024b] and the computational scalability for large-scale systems is investigated in [Stomberg2024c]. Specifically, this article presents two contributions: First, we derive novel dSQP convergence guarantees when the number of inner iterations is fixed instead of relying on an inexact Newton type stopping criterion as in [Stomberg2022a]. Second, we combine the linear convergence of dSQP with the RTI stability guarantees from [Zanelli2021] to derive the local exponential stability of the system-optimizer dynamics in closed loop. To the best of our knowledge, we are the first to study the system-optimizer convergence in DMPC and provide the respective ADMM convergence guarantees.
@@
The article is structured as follows: Section II states the control objective and presents the OCP. Section III recalls RTI stability for centralized NMPC. Section IV explains the bi-level dSQP scheme and derives new q-linear convergence guarantees when the number of inner iterations per outer iterations is fixed. Section V presents the stability of the distributed RTI scheme. Section VI analyzes numericals results for coupled inverted pendulums.
@@
Notation: Given a matrix $A$ and an integer $j$, $[A]_{j}$ denotes the $j$th row of $A$. For an index set $\mathcal{A}$, $[A]_{\mathcal{A}}$ denotes the matrix consisting of rows $[A]_{j}$ for all $j\in\mathcal{A}$. Likewise, $[a]_{j}$ is the $j$th component of vector $a$ and $a_{\mathcal{A}}$ is the vector of components $[a]_{j}$ for all $j\in\mathcal{A}$. The concatenation of vectors $x$ and $y$ into a column vector is $(x,y)$. Given scalars $a_{1},\dots,a_{n}$, $A=\mathrm{diag}(a_{1},\dots,a_{n})\in\mathbb{R}^{n\times n}$ is the diagonal matrix where $[A]_{ii}=a_{i}$. Likewise, given matrices $A_{1},\dots,A_{S}$, $C=\mathrm{diag}(A_{1},\dots,A_{S})$ is a block diagonal matrix with block $[C]_{ii}=A_{i}$. The Euclidean norm of a vector $a\in\mathbb{R}^{n}$ is denoted by $\|a\|\doteq\sqrt{a^{\top}a}$. The spectral norm of a matrix $A\in\mathbb{R}^{n\times m}$ is denoted by $\|A\|\doteq\sigma_{\max}(A)$, the largest singular value of $A$. The closed $\varepsilon$ neighborhood around a point $x^{\star}\in\mathbb{R}^{n}$ is denoted as $\mathcal{B}(x^{\star},\varepsilon)$, i.e., $\mathcal{B}(x^{\star},\varepsilon)\doteq\{x\in\mathbb{R}^{n}|\|x-x^{\star}\|\leq\varepsilon\}$. We denote the natural numbers by $\mathbb{N}$, the natural numbers extended by zero by $\mathbb{N}_{0}\doteq\{0\}\cup\mathbb{N}$, the set of integers by $\mathbb{I}$, the set of integers in the range from 00 to $N$ by $\mathbb{I}_{[0,N]}$, and the Minkowski sum of sets $A$ and $B$ as $A\oplus B$. Given $a\in\mathbb{R}$, $b=\lceil a\rceil$ is the nearest integer $b\geq a$.
II Problem Statement
@@
[…]
$\bar{x}_{i}(\tau+1)$ $=f_{i}^{\delta}(\bar{x}_{i}(\tau),\bar{u}_{i}(\tau),\bar{x}_{\mathcal{N}_{i}^{\text{in}}}(\tau)),\hskip 11.09654pt\forall\tau\in\mathbb{I}_{[0,N-1]},$ (1b)
$\bar{x}_{i}(0)$ $=x_{i}(t),$ (1c)
$\bar{x}_{i}(\tau)$ $\in\mathbb{X}_{i},\,\forall\tau\in\mathbb{I}_{[0,N]},\,\bar{u}_{i}(\tau)\in\mathbb{U}_{i},\,\forall\tau\in\mathbb{I}_{[0,N-1]},$ (1d)
$(\bar{x}_{i}(\tau),$ $\bar{x}_{j}(\tau))\in\mathbb{X}_{ij},\hskip 28.45274pt\forall j\in\mathcal{N}_{i}^{\text{in}},\;\forall\tau\in\mathbb{I}_{[0,N]},$ (1e)
@@
The state and input of subsystem $i$ are denoted by $x_{i}\in\mathbb{R}^{n_{x_{i}}}$ and $u_{i}\in\mathbb{R}^{n_{u_{i}}}$, respectively. To distinguish closed-loop and open-loop trajectories, we denote predicted states and inputs with a superscript $\bar{\cdot}$. The decision variables of OCP (1) are the predicted state trajectories $\bar{\boldsymbol{x}}_{i}$ and input trajectories $\bar{\boldsymbol{u}}_{i}$ over the horizon $N$. Define $n_{i}^{\text{in}}\doteq\sum_{j\in\mathcal{N}_{i}^{\text{in}}}n_{x_{j}}$. We stack the states of in-neighbors of subsystem $i$ in alphabetical order in the vector ${x}_{\mathcal{N}_{i}^{\text{in}}}\in\mathbb{R}^{n_{i}^{\text{in}}}$. The objective (1a) consists of individual stage costs $\ell_{i}:\mathbb{R}^{n_{x_{i}}}\times\mathbb{R}^{n_{u_{i}}}\times\mathbb{R}^{n_{i}^{\text{in}}}\rightarrow\mathbb{R}$, terminal penalties $V_{\text{f},i}:\mathbb{R}^{n_{x_{i}}}\rightarrow\mathbb{R}$, and a scaling factor $\beta\geq 1$. For all $i\in\mathcal{S}$, $f_{i}^{\delta}:\mathbb{R}^{n_{x_{i}}}\times\mathbb{R}^{n_{u_{i}}}\times\mathbb{R}^{n_{i}^{\text{in}}}\rightarrow\mathbb{R}^{n_{x_{i}}}$ denotes the discrete-time dynamics with control sampling interval $\delta>0$. The states and inputs are constrained to the sets $\mathbb{X}_{i}\subseteq\mathbb{R}^{n_{x_{i}}}$, $\mathbb{U}_{i}\subseteq\mathbb{R}^{n_{u_{i}}}$. Each set $\mathbb{X}_{ij}\subseteq\mathbb{R}^{n_{x_{i}}}\times\mathbb{R}^{n_{x_{j}}}$ couples two neighbors $i,j\in\mathcal{S}$. We do not enforce additional terminal constraints to reduce the computational burden [Limon2006].
@@
$\bar{x}(\tau+1)$ $=f^{\delta}(\bar{x}(\tau),\bar{u}(\tau)),\hskip 43.24817pt\forall\tau\in\mathbb{I}_{[0,N-1]},$ (2b)
@@
The centralized system state and input are $x=(x_{1},\dots,x_{S})\in\mathbb{R}^{n_{x}}$ and $u=(u_{1},\dots,u_{S})\in\mathbb{R}^{n_{u}}$, respectively. The centralized discrete-time dynamics $f^{\delta}:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{u}}\rightarrow\mathbb{R}^{n_{x}}$ are obtained by sampling the corresponding continuous-time dynamics $f^{\mathrm{c}}:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{u}}\rightarrow\mathbb{R}^{n_{x}}$ with piecewise constant input signals at $\delta>0$. The partitioning of the centralized system into subsystems affects the performance of DMPC and is studied in [Chanfreut2021].
The elements of OCP (2) are comprised of the components of (1). Considering the entire network as a single system of high state dimension serves as a conceptual means in the stability analysis and allows us to draw upon existing RTI stability guarantees. Notice that the numerical scheme to be proposed subsequently is decentralized, because we solve OCP (1) via dSQP [Stomberg2022a].
III Centralized Real-Time Iterations
RTI schemes are designed to ensure the nominal closed-loop properties, if only few optimizer iterations are taken in each control step to compute a control input for system (2b). This section recalls stability guarantees that also hold when inequality constraints are present in the OCP [Zanelli2021].
To this end, we define the NMPC control law $\kappa_{\mathrm{c}}:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}^{n_{u}}$ as the map from the initial state $x(t)$ in (2c) to the first part $\bar{u}^{\star}(0)$ of the globally optimal input trajectory. We then make the following assumption on the value function $V:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}$, which can be met by appropriate OCP design, cf. [Limon2006]. OCP designs specific to DMPC are discussed in [Dunbar2007, Conte2016, Darivianakis2019]. Define the level set $\mathbb{X}_{\bar{V}}\doteq\{x\in\mathbb{R}^{n_{x}}|V(x)\leq\bar{V}\}$ for a constant $\bar{V}>0$.
Assumption 1 (Value function requirements [Zanelli2021])
The value function $V:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}$ of OCP (2) is continuous and there exist positive constants $a_{1}$, $a_{2}$, $a_{3}$, $\delta_{0}$, and $\bar{V}$ such that, for all $x\in\mathbb{X}_{\bar{V}}$ and for all sampling intervals $\delta\leq\delta_{0}$,

### v3

@@ Abstract
{IEEEkeywords}
Nonlinear model predictive control, real-time iterations, decentralized optimization, distributed control, alternating direction method of multipliers
1 Introduction
@@
Distributed control concerns the operation and control of cyber-physical systems, e.g., energy systems [1] or robot formations [2]. Coupling in systems of systems can occur in the dynamics, via constraints, or through a common objective. A key challenge for the design of optimization-based schemes for interconnected and cyber-physical systems is to reconcile cooperation with the computational burden, i.e., real-time feasibility is a must in applications. On the far end of the spectrum, decentralized control schemes do not allow for cooperation but also do not require communication between subsystems [3]. The opposite is centralized MPC, where the network of subsystems is considered as one, large-scale, dynamical system which is controlled by a single controller.
@@
Model Predictive Control (MPC), also known as receding-horizon optimal control, has seen tremendous industrial success catalyzed by the development of numerical schemes tailored to the dynamics and to the problem formulation. Of particular importance for the implementation of centralized Nonlinear MPC (NMPC) are Real-Time Iteration (RTI) schemes [4, 5, 6, 7, 8, 9]. Instead of solving the Optimal Control Problem (OCP) to full accuracy in each control step, an RTI scheme applies only one or a few iterations of an optimization method.
@@
On the other hand, Distributed MPC (DMPC) decomposes the numerical optimization among the subsystems [10, 11]. DMPC design proceeds mainly along two dimensions: (i) the OCP formulation and (ii) the implementation of an optimization algorithm with the desired degree of decomposition or decentralization.
@@
With respect to (i), one may design either an individual OCP for each subsystem [12, 13, 14], or a centralized OCP for the whole cyber-physical system [15, 16]. Schemes with individual OCPs for the subsystems require few communication rounds per control step, enjoy small communication footprints, and allow for fast sampling rates [12, 13, 14]. However, they only allow for limited cooperation. Distributed approaches built upon centralized OCPs generally require multiple communication rounds per control step, but also allow for cooperation as the control tasks of all subsystems are encoded in the centralized OCP. We therefore refer to the latter approach as cooperative DMPC in this article.
@@
With respect to (ii), numerous optimization algorithms have been proposed to solve centralized OCPs online [17, 18, 19, 20]. A distinction can be made between distributed and decentralized optimization methods. Distributed optimization splits most computations between the subsystems and there exists a central entity which coordinates the subsystems [21]. Decentralized optimization only requires neighbor-to-neighbor communication without a coordinator [22].
@@
Ultimately, cooperative DMPC aims to combine the high performance of centralized MPC with the favorable communication structure of decentralized optimization. Real-time requirements dictate that only a finite number of optimizer iterations can be executed in each control step, which limits performance and must be addressed in the stability analysis. While RTI schemes have been pivotal in bringing NMPC to industrial applications, distributed counterparts for similar results are not available. For linear systems, DMPC-specific OCP designs are presented in [15, 23] and stability under inexact optimization is analyzed in [24, 25]. For nonlinear systems, cooperative DMPC schemes with stability guarantees are presented in [17, 16]. In both articles, a centralized OCP with a terminal penalty is formulated such that the OCP value function serves as a candidate Lyapunov function for the closed-loop system. Stability is ensured in two different ways: Either a feasible-side convergent optimization method is employed and stability follows from standard arguments [17]. The drawback of this approach is that the optimization method requires a feasible initialization, which is difficult to implement in practice, and that each subsystem needs access to the dynamics of all subsystems in the network. The second option for guaranteeing stability is to solve the OCP with the Alternating Direction Method of Multipliers (ADMM) until a tailored ADMM stopping criterion is met [16]. This approach presumes that ADMM converges linearly, i.e., sufficiently fast, to the OCP minimizer. However, such ADMM convergence guarantees for problems with non-convex constraint sets are, to the best of the authors’ knowledge, yet unavailable. Moreover, we note that the existing ADMM convergence guarantees for non-convex constraint sets [26, 27] so far do not allow for decentralized implementations.
@@
Consequently, the existing stability guarantees of cooperative DMPC for coupled nonlinear systems either require feasible initialization and global model knowledge in all subsystems or they rely on rather strong assumptions on the achieved optimizer convergence. To overcome these limitations, we propose a novel decentralized RTI scheme for distributed NMPC. Our scheme builds on a bi-level decentralized Sequential Quadratic Programming (dSQP) scheme [28]. On the outer level, the method uses an inequality-constrained SQP scheme, which leads to partially separable convex Quadratic Programs (QPs) to be solved in each SQP step. On the inner level, these QPs are solved with ADMM which is guaranteed to converge and can be implemented in decentralized fashion, i.e., it does not require a central coordinator. Our previous conference paper [28] presents an earlier dSQP version including a stopping criterion for ADMM, but it does not consider real-time control applications and stability analysis. The idea of executing only few iterations of a tailored optimization method to enable distributed NMPC in real-time is also used in [19]. Therein, an augmented Lagrangian-based decomposition scheme is proposed for non-convex OCPs and suboptimality bounds for the optimizer solutions are derived, if changes in the system state between subsequent NMPC steps are small. While stability of the dynamical system is not formally discussed in [19], the approach shares commonalities to the scheme developed in this article and we later give a more detailed comparison in Remark 16 in Section 5.
@@
In the present paper, we explore the theoretical foundation of decentralized real-time iterations for DMPC via dSQP. The real-time feasibility of our approach has successfully been validated in experiments with mobile robots and on embedded hardware [29, 30] and the computational scalability for large-scale systems is investigated in [31]. Specifically, this article presents two contributions: First, we derive novel dSQP convergence guarantees when the number of inner iterations is fixed instead of relying on an inexact Newton type stopping criterion as in [28]. Second, we combine the linear convergence of dSQP with the RTI stability guarantees from [7] to derive the local exponential stability of the system-optimizer dynamics in closed loop. To the best of our knowledge, we are the first to study the system-optimizer convergence in DMPC and provide the respective ADMM convergence guarantees.
@@
The article is structured as follows: Section 2 states the control objective and presents the OCP. Section 3 recalls RTI stability for centralized NMPC. Section 4 explains the bi-level dSQP scheme and derives new q-linear convergence guarantees when the number of inner iterations per outer iteration is fixed. Section 5 presents the stability of the distributed RTI scheme. Section 6 analyzes numericals results for coupled inverted pendulums.
@@
Notation: Given a matrix $A$ and an integer $j$, $A]_{j}$ denotes the $j$th row of $A$. For an index set $\mathcal{A}$, $A]_{\mathcal{A}}$ denotes the matrix consisting of rows $A]_{j}$ for all $j\in\mathcal{A}$. Likewise, $a]_{j}$ is the $j$th component of vector $a$ and $a_{\mathcal{A}}$ is the vector of components $a]_{j}$ for all $j\in\mathcal{A}$. The concatenation of vectors $x$ and $y$ into a column vector is $(x,y)$. Given scalars $a_{1},\dots,a_{n}$, $A=\mathrm{diag}(a_{1},\dots,a_{n})\in\mathbb{R}^{n\times n}$ is the diagonal matrix where $A]_{ii}=a_{i}$. Likewise, given matrices $A_{1},\dots,A_{S}$, $C=\mathrm{diag}(A_{1},\dots,A_{S})$ is a block diagonal matrix with block $C]_{ii}=A_{i}$. The Euclidean norm of a vector $a\in\mathbb{R}^{n}$ is denoted by $\|a\|\doteq\sqrt{a^{\top}a}$. The spectral norm of a matrix $A\in\mathbb{R}^{n\times m}$ is denoted by $\|A\|\doteq\sigma_{\max}(A)$, the largest singular value of $A$. The closed $\varepsilon$ neighborhood around a point $x^{\star}\in\mathbb{R}^{n}$ is denoted as $\mathcal{B}(x^{\star},\varepsilon)$, i.e., $\mathcal{B}(x^{\star},\varepsilon)\doteq\{x\in\mathbb{R}^{n}|\|x-x^{\star}\|\leq\varepsilon\}$. We denote the natural numbers by $\mathbb{N}$, the natural numbers extended by zero by $\mathbb{N}_{0}\doteq\{0\}\cup\mathbb{N}$, the set of integers by $\mathbb{I}$, the set of integers in the range from 0 to $N$ by $\mathbb{I}_{[0,N]}$, and the Minkowski sum of sets $A$ and $B$ as $A\oplus B$. Given $a\in\mathbb{R}$, $b=\lceil a\rceil$ is the nearest integer $b\geq a$.
2 Problem Statement
@@
subject $\text{ to }\text{for all }i\in\mathcal{S}$
$\hskip-8.53581pt\bar{x}_{i}(\tau\hskip-1.42262pt+\hskip-1.42262pt1)$ $=f_{i}^{\delta}(\bar{x}_{i}(\tau),\bar{u}_{i}(\tau),\bar{x}_{\mathcal{N}_{i}^{\text{in}}}(\tau)),\hskip 15.649pt\forall\tau\in\mathbb{I}_{[0,N-1]},$ (1b)
$\hskip-8.53581pt\bar{x}_{i}(0)$ $=x_{i}(t),$ (1c)
$\hskip-8.53581pt\bar{x}_{i}(\tau)$ $\in\mathbb{X}_{i},\;\forall\tau\in\mathbb{I}_{[0,N]},\;\,\bar{u}_{i}(\tau)\in\mathbb{U}_{i},\;\forall\tau\in\mathbb{I}_{[0,N-1]},$ (1d)
$\hskip-8.53581pt(\bar{x}_{i}(\tau),$ $\bar{x}_{j}(\tau))\in\mathbb{X}_{ij},\hskip 33.57425pt\forall j\in\mathcal{N}_{i}^{\text{in}},\;\forall\tau\in\mathbb{I}_{[0,N]},$ (1e)
@@
The state and input of subsystem $i$ are denoted by $x_{i}\in\mathbb{R}^{n_{x_{i}}}$ and $u_{i}\in\mathbb{R}^{n_{u_{i}}}$, respectively. To distinguish closed-loop and open-loop trajectories, we denote predicted states and inputs with a superscript $\bar{\cdot}$. The decision variables of OCP (1) are the predicted state trajectories $\bar{\boldsymbol{x}}_{i}$ and input trajectories $\bar{\boldsymbol{u}}_{i}$ over the horizon $N$. Define $n_{i}^{\text{in}}\doteq\sum_{j\in\mathcal{N}_{i}^{\text{in}}}n_{x_{j}}$. We stack the states of in-neighbors of subsystem $i$ in alphabetical order in the vector ${x}_{\mathcal{N}_{i}^{\text{in}}}\in\mathbb{R}^{n_{i}^{\text{in}}}$. The objective (1a) consists of individual stage costs $\ell_{i}:\mathbb{R}^{n_{x_{i}}}\times\mathbb{R}^{n_{u_{i}}}\times\mathbb{R}^{n_{i}^{\text{in}}}\rightarrow\mathbb{R}$, terminal penalties $V_{\text{f},i}:\mathbb{R}^{n_{x_{i}}}\rightarrow\mathbb{R}$, and a scaling factor $\beta\geq 1$. For all $i\in\mathcal{S}$, $f_{i}^{\delta}:\mathbb{R}^{n_{x_{i}}}\times\mathbb{R}^{n_{u_{i}}}\times\mathbb{R}^{n_{i}^{\text{in}}}\rightarrow\mathbb{R}^{n_{x_{i}}}$ denotes the discrete-time dynamics with control sampling interval $\delta>0$. The states and inputs are constrained to the closed sets $\mathbb{X}_{i}\subseteq\mathbb{R}^{n_{x_{i}}}$, $\mathbb{U}_{i}\subseteq\mathbb{R}^{n_{u_{i}}}$. Each closed set $\mathbb{X}_{ij}\subseteq\mathbb{R}^{n_{x_{i}}}\times\mathbb{R}^{n_{x_{j}}}$ couples two neighbors $i,j\in\mathcal{S}$. We do not enforce additional terminal constraints to reduce the computational burden [32].
@@
$\bar{x}(\tau+1)$ $=f^{\delta}(\bar{x}(\tau),\bar{u}(\tau)),\hskip 49.50777pt\forall\tau\in\mathbb{I}_{[0,N-1]},$ (2b)
@@
The centralized system state and input are $x=(x_{1},\dots,x_{S})\in\mathbb{R}^{n_{x}}$ and $u=(u_{1},\dots,u_{S})\in\mathbb{R}^{n_{u}}$, respectively. The centralized discrete-time dynamics $f^{\delta}:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{u}}\rightarrow\mathbb{R}^{n_{x}}$ are obtained by sampling the corresponding continuous-time dynamics $f^{\mathrm{c}}:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{u}}\rightarrow\mathbb{R}^{n_{x}}$ with piecewise constant input signals at $\delta>0$. The partitioning of the centralized system into subsystems affects the performance of DMPC and is studied in [33].
The elements of OCP (2) are comprised of the components of (1). Considering the entire network as a single system of high state dimension serves as a conceptual means in the stability analysis and allows us to draw upon existing RTI stability guarantees. Notice that the numerical scheme to be proposed subsequently is decentralized, because we solve OCP (1) via dSQP [28].
3 Centralized Real-Time Iterations
RTI schemes are designed to ensure the nominal closed-loop properties if only few optimizer iterations are taken in each control step to compute a control input for system (2b). This section recalls stability guarantees that also hold when inequality constraints are present in the OCP [7].
Let $\mathbb{X}_{0}\subseteq\mathbb{R}^{n_{x}}$ be a closed set with nonempty interior and let OCP (2) be feasible for all initial states $x(t)\in\mathbb{X}_{0}$. We define the NMPC control law $\kappa_{\mathrm{c}}:\mathbb{X}_{0}\rightarrow\mathbb{R}^{n_{u}}$ as the map from the initial state $x(t)$ in (2c) to the first part $\bar{u}^{\star}(0)$ of a globally optimal input trajectory. We then make the following assumption on the value function $V:\mathbb{X}_{0}\rightarrow\mathbb{R}$, which can be met by appropriate OCP design, cf. [32]. OCP designs specific to DMPC are discussed in [12, 15, 23].
Assumption 1 (Value function requirements [7])
The value function $V:\mathbb{X}_{0}\rightarrow\mathbb{R}$ of OCP (2) is continuous and there exist positive constants $a_{1}$, $a_{2}$, $a_{3}$, and $\bar{V}$ such that, for all $x\in\mathbb{X}_{\bar{V}}\doteq\{x\in\mathbb{R}^{n_{x}}\,|\,V(x)\leq\bar{V}\}$,